The length of a rectangle is 5 feet longer than its width. If its area is 36 feet, what are its dimensions?
Answer by algebrahouse.com(1659)

x = width
x + 5 = length {length is 5 longer than width}

Area of a rectangle = width x length

x(x + 5) = 36 {substituted width and length into area formula}
x² + 5x = 36 {used distributive property}
x² + 5x - 36 = 0 {subtracted 36 from each side}
(x + 9)(x - 4) = 0 {factored into two binomials}
x + 9 = 0 or x - 4 = 0 {set each factor equal to 0}
x = -9 or x = 4 {solved each equation for x}

x = 4 {width cannot be negative}
x + 5 = 9 {substituted 4, in for x, into x + 5}

width = 4 ft
length = 9 ft

For more help from me, visit: www.algebrahouse.com